OpenStudy (anonymous):

Please help! Solve for x: x=log_4 5. An intermediate step is?

OpenStudy (anonymous):

Is this the entire question? Cause I know how to work this question, the question just isn't complete.

OpenStudy (sasogeek):

log_4 (5) is a numerical value, and x = log_4 (5) , there's nothing to solve, unless you want to find the exact value, use a calculator :)

OpenStudy (kropot72):

\[x=\log_{4} 5\] \[4^{x}=5\] \[x \ln 4=\ln 5\] \[x=\frac{\ln 5}{\ln 4}\]

OpenStudy (anonymous):

@Johnjakile1998 Yea, it's the complete question. It's just asking for an intermediate step when solving :/ Thank you @sasogeek and @kropot72 :)

OpenStudy (anonymous):

So, do you still need help?

OpenStudy (kropot72):

You're welcome :)

OpenStudy (dumbcow):

what @kropot72 did was show how to get x in terms of natural log most calculators only have natural log or log base 10 the short version is known as change of base formula \[\log_{b}A = \frac{\ln A}{\ln b} = \frac{\log A}{\log b} \]

OpenStudy (sasogeek):

the natural log is just log to the base e. truth is, log to any base will do, but it's best if we stick to the standards, just saying :) and here's what i'm saying :P \(\huge log_bA=\frac{log_{10}A}{log_{10}b}=\frac{log_{5}A}{log_{5}b}=\frac{lnA}{lnb}=\frac{log_{23}A}{log_{23}b} \) and so on and so forth :)

OpenStudy (anonymous):

Thank you @dumbcow Sorry, OS keeps lagging on me :(

OpenStudy (kropot72):

So we can say \[\log_{b} A=\frac {\log_{} A}{\log_{} b}\]